OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2.

OpenStudy (anonymous):

help will give medal

OpenStudy (anonymous):

have you done sequences before?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

ok, so we have the formula: f(n) = f(1) + f(2) + f(n - 1) every where you see a f(1), replace it with a 2 becuase f(1)=2 everywhere you see a f(2), replace it with a 4 becuase f(2)=4 now, what does the formula look like now?

OpenStudy (anonymous):

confusing still

OpenStudy (anonymous):

thanks for being honest! I'll do the first part so you can see... the formula: f(n) = f(1) + f(2) + f(n - 1) every where you see a f(1), replace it with a 2 becuase f(1)=2 So it becomes f(n) = 2 + f(2) + f(n - 1) you try the other part, "everywhere you see a f(2), replace it with a 4 becuase f(2)=4"

OpenStudy (anonymous):

so whats the answer

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=C7tE26CDI2M

OpenStudy (anonymous):

whats the answer though

OpenStudy (anonymous):

you'll have to find someone else to help you if you are looking for answers, cya!

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

To find f(5) you just sustitute 5 in the formula.